### v1

@@ Abstract
This paper uses the flow cytometry dataset provided by our University, which will be available in our GitHub benchmark’s repository along with our code, after the reviewing process.
@@ 1 Introduction
We endeavor to offer the first public benchmark for Flow Cytometry data, providing automated solutions for hematological relevant tasks, such as multi-class single-cell classification. It is noteworthy that this is the first available benchmark for Flow Cytometry data, equipped with baseline models and advanced deep learning methods. This benchmark aims to provide a standardized platform for evaluating the performance of various algorithms and models, facilitating the development of novel, more efficient, and accurate diagnostic strategies.
@@ 11 CD45-KO Cluster of Differentiation 45 - Antigen
Overall, the data repository includes files for a total volume of around 1.3 GB. The data is distributed in the convenient GitHub repository [^4] Our anonymous GitHub.
@@ 4.1.1 Baselines
To use these graph learning methods, FlowCyt provides the option to convert tabular data to graph-structured data that can be fed into the various proposed GNNs. We achieved this by treating each cell as a node in the graph space, with $\mathcal{N}_{i}$ representing the set of $k$-nearest neighbors for node $i$ (where $k=7$) using the $\text{L}_{2}$ metric in the feature space. For each node, we built all possible pairs of its neighbors and added them to the edge set of the graph, resulting in a fully connected graph for each patient-graph instance. After experimenting with different values, we decided to use $k=7$ to strike a balance between making the graph too sparse (which could miss important connections) and too dense (which could include irrelevant connections). This approach enabled us to capture both the local structure (individual neighbors) and higher-order relationships (combinations of neighbors) in the data, helping us identify the most relevant relationships without adding noise.
@@ 4.1.2 Inductive Learning
Motivated by the best results produced in Section 4.1.2, we further proved the effectiveness of the models by using the transductive learning methodology outlined in Yang et al. (2016). During the training phase, the models were given access to the feature vectors of all nodes. Our primary goal was to classify the status of cells, distinguishing between the five different classes, while also evaluating the reliability of the classification performances in a semi-supervised setup utilizing up to one million cells per patient.
To achieve this, we conducted five different seed iterations using our models. We randomly assigned 10% of the nodes for validation and 10% for testing while keeping the classes’ ratio balanced, the same as in the full dataset. Since our model is strongly imbalanced, we used a weighted negative log-likelihood loss function. Then we randomly masked the labels of 50% of the training, validation, and test set. We can therefore think of our problem as having a large graph where half of the graph is unlabeled, and our goal is to predict the labels based on the ones that we have.
@@ 4.2.3 Populations on TL
In this section, we discuss the results obtained for the sub-population on TL set-up (see Section 3.1.1). We have evaluated our metrics and found that the F1 score is around $\left\{0.94\pm 0.01\right\}$ for GAT, GCN, and SAGE, which confirms the robustness of graph models and graph encoded data even for semi-supervised approach with large graphs. The results for the employed models have been obtained using those hyperparameters to fit the constraints of 1 GPU with 11GB memory, as shown in Table LABEL:tab:hyper_trans of Appendix A. To highlight the strength of graph models, Figure LABEL:fig:tSNE23 shows the t-SNE (Van der Maaten and Hinton, 2008) embeddings visualization of GAT for one random patient (patient 23 has been randomly picked). Despite the difficulties due to the mask setup that is typical of transductive learning, we can see how the graph model has clustered the five classes into its final embedding layer. To view the results for the sub-population under the TL framework, please refer to our code repository.
@@
Figure 3: t-SNE projection for patient one random patient, for the transductive learning task.
@@ 5 Discussion of FlowCyt
The feasibility of developing automated solutions to assist cytometrists has been demonstrated through multi-class classification experiments throughout Section 4. The effectiveness of graph neural networks in modeling spatial patterns and inter-cell dependencies via graph encoding has been highlighted by the results obtained. This benchmark represents the first public flow cytometry benchmark that empowers rigorous assessment of novel methodologies. It will facilitate research that pushes the boundaries and clinical translation of automated single-cell analysis to supplement traditional manual examination.
Through collaborative efforts, this versatile dataset may uncover additional tasks and biological insights, thus contributing to advancing the field of hematological cell populations (Bini et al., 2024).
@@ 5.1 Future Research
This benchmark is a solid foundation for expanding research in flow cytometry through diverse data and advanced analytical tasks. To better represent real clinical scenarios and improve model robustness, it is important to increase our patient cohort and include patients with hematologic/immunologic diseases, such as leukemias. Obtaining longitudinal samples that track patients over treatment courses could also reveal informative dynamics and could help to understand how cell populations evolve in response to therapies.

### v2

@@ Abstract
This paper uses the flow cytometry dataset provided by our University Hospital, which is available in our GitHub benchmark’s repository along with our code [^1] FlowCyt Webpage.
@@ 1 Introduction
We endeavor to offer the first public benchmark for flow cytometry data, providing automated solutions for hematological relevant tasks, such as multi-class single-cell classification. It is noteworthy that this is the first available benchmark for flow cytometry data, equipped with ground truth, baseline models and advanced deep learning methods. This benchmark aims to provide a standardized platform for evaluating the performance of various algorithms and models, facilitating the development of novel, more efficient, and accurate diagnostic strategies.
@@ 11 CD45-KO Cluster of Differentiation 45 - Antigen
While we acknowledge that our study utilizes a dataset with 12 markers, this number is well within the range of typical flow cytometry panels used in clinical practice. Many previous studies have successfully demonstrated the utility of machine learning techniques on similar or even lower dimensional flow data. For instance, Arvaniti and Claassen (2017) applied representation learning on a 10-marker dataset to detect rare disease populations. Bruggner et al. (2014)identified stratifying signatures across 13 surface markers, while Finak et al. (2009) proposed a mixture model-based approach for cell population identification on a datasets with 8-12 markers, showcasing the effectiveness of their method in accurately identifying cell populations across diverse marker panels. Moreover, the key challenge in FC analysis does not solely stem from high dimensionality but rather from the inherent heterogeneity, complexity, and rarity of cell populations of interest amidst a majority of irrelevant cells. Our dataset, comprising over 1 million cells per patient, with 5 and 6 distinct and unbalanced cell types, accurately reflects these real-world challenges faced by cytometrists.
@@
Overall, the data repository includes files for a total volume of around 1.3 GB. The data is distributed in our convenient group repository [^5] FlowCyt Webpage.
@@ 4.1.1 Baselines
To use these graph learning methods, FlowCyt provides the option to convert tabular data to graph-structured data that can be fed into the various proposed GNNs. We achieved this by treating each cell as a node in the graph space, with $\mathcal{N}_{i}$ representing the set of $k$-nearest neighbors for node $i$ (where $k=7$) using the $\text{L}_{2}$ metric in the feature space. For each node, we built all possible pairs of its neighbors and added them to the edge set of the graph, resulting in a fully connected graph for each patient-graph instance. After experimenting with different values, we decided to use $k=7$ to strike a balance between making the graph too sparse (which could miss important connections) and too dense (which could include irrelevant connections). This approach enabled us to capture both the local structure (individual neighbors) and higher-order relationships (combinations of neighbors) in the data, helping us identify the most relevant relationships without adding noise. Moreover, constructing the graph via the $k$-nearest neighbor is intrinsically local and does not force us to make any assumptions about the distribution of the data, making it a good choice for datasets like ours where the full distribution of the data could be unknown.
@@ 4.1.2 Inductive Learning
Motivated by the best results produced in Section 4.1.2, we further proved the effectiveness of the models by using the transductive learning methodology outlined in Yang et al. (2016), where during the training phase the models were given access to the feature vectors of all nodes. While the concept of transductive learning has been explored in various domains, its application to flow cytometry data is relatively novel, and there is little literature explicitly discussing its implications in this field. Hu et al. (2020) and Zhang et al. (2019) proposed a robust and interpretable end-to-end deep learning model for cytometry data and scRNA-Seq analysis, incorporating elements of self-supervised and semi-supervised learning. They discuss the advantages of leveraging unlabeled data to improve model generalization and robustness, which is crucial in the context of heterogeneous and high-dimensional cytometry data to reduce the need for extensive manual gating and annotation.
In FlowCyt, our primary goal was to classify the status of cells, distinguishing between the five different classes, while also evaluating the reliability of the classification performances in a semi-supervised setup utilizing up to one million cells per patient. To achieve this, we conducted five different seed iterations using our models. We randomly assigned 10% of the nodes for validation and 10% for testing while keeping the classes’ ratio balanced, the same as in the full dataset. Since our model is strongly imbalanced, we used a weighted negative log-likelihood loss function. Then we randomly masked the labels of 50% of the training, validation, and test set. We can therefore think of our problem as having a large graph where half of the graph is unlabeled, and our goal is to predict the labels based on the ones that we have.
@@ 4.2.3 Populations on TL
In this section, we discuss the results obtained for the sub-population on TL set-up (see Section 3.1.1). We have evaluated our metrics and found that the F1 score is around $\left\{0.94\pm 0.01\right\}$ for GAT, GCN, and SAGE, which confirms the robustness of graph models and graph encoded data even for semi-supervised approach with large graphs. The results for the employed models have been obtained using those hyperparameters to fit the constraints of 1 GPU with 11GB memory (RTX 2080 Ti), as shown in Table LABEL:tab:hyper_trans of Appendix A. To highlight the strength of graph models, Figure LABEL:fig:tSNE23 shows the t-SNE (Van der Maaten and Hinton, 2008) embeddings visualization of GAT for one random patient (patient 23 has been randomly picked). Despite the difficulties due to the mask setup that is typical of transductive learning, we can see how the graph model has clustered the five classes into its final embedding layer. To view the results for the sub-population under the TL framework, please refer to our code repository.
@@
Figure 3: t-SNE projection for one random patient, for the transductive learning task.
@@ 5 Discussion of FlowCyt
The feasibility of developing automated solutions to assist cytometrists has been demonstrated through multi-class classification experiments throughout Section 4. The effectiveness of graph neural networks in modeling spatial patterns and inter-cell dependencies via graph encoding has been highlighted by the results obtained. This benchmark represents the first public flow cytometry benchmark that empowers rigorous assessment of novel methodologies. It will facilitate research that pushes the boundaries and clinical translation of automated single-cell analysis to supplement traditional manual examination. Through collaborative efforts, this versatile dataset may uncover additional tasks and biological insights, thus contributing to advancing the field of hematological cell populations (Bini et al., 2024).
@@ 5.1 Future Research
This benchmark is a solid foundation for expanding research in flow cytometry through diverse data and advanced analytical tasks. To better represent real clinical scenarios and improve model robustness, it is important to increase our patient cohort and include patients with hematologic/immunologic diseases, such as leukemias (e.g. Acute Myeloid Leukemia-AML). Obtaining longitudinal samples that track patients over treatment courses could also reveal informative dynamics and could help to understand how cell populations evolve in response to therapies.